You’re traveling 45 km at 30o north east, then 98 km at 50o north west. What is your resultant displacement?


1
Expert's answer
2021-11-14T17:08:59-0500

Let's first find the "x"- and "y"- components of the resultant displacement:


"d_x=45\ km\times cos30^{\circ}+98\ km\times cos(180^{\circ}-50^{\circ})=-24\ km,""d_y=45\ km\times sin30^{\circ}+98\ km\times sin(180^{\circ}-50^{\circ})=97.6\ km."

We can find the magnitude of the resultant displacement from the Pythagorean theorem:


"d=\sqrt{d_x^2+d_y^2}=\sqrt{(-24\ km)^2+(97.6\ km)^2}=100.5\ km."

We can find the direction of the resultant displacement from the geometry:


"\theta=tan^{-1}(\dfrac{d_y}{d_x}),""\theta=tan^{-1}(\dfrac{97.6\ km}{-24\ km})=76.2^{\circ} N\ of\ W."